### v2

# CAVE: Controllable Authorship Verification Explanations

Abstract
Authorship Verification (Av) (do two documents have the same author?) is essential in many sensitive real-life applications. Av is often used in proprietary domains that require a private, offline model, making SOTA online models like ChatGPT undesirable. Current offline models however have lower downstream utility due to low accuracy/scalability (eg: traditional stylometry Av systems) and lack of accessible post-hoc explanations. In this work, we take the first step to address the above challenges with our trained, offline LLaMa-3-8B model Cave (Controllable Authorship Verification Explanations): Cave generates free-text Av explanations that are controlled to be (1) structured (can be decomposed into sub-explanations in terms of relevant linguistic features), and (2) easily verified for explanation-label consistency (via intermediate labels in sub-explanations). We first engineer a prompt that can generate silver training data from a SOTA teacher model in the desired Cave output format. We then filter and distill this data into a pretrained LLaMa-3-8B, our carefully selected student model. Results on three difficult Av datasets IMDb62, Blog-Auth, and FanFiction show that Cave generates high quality explanations (as measured by automatic and human evaluation) as well as competitive task accuracies111github.com/INK-USC/Controllable-AV-Explanations .
CAVE: Controllable Authorship Verification Explanations
Sahana $\varheartsuit\varheartsuit{}^{\varheartsuit}$, Kartik $\varheartsuit\varheartsuit{}^{\varheartsuit}$, Elizabeth Boschee, Xiang $\varheartsuit\varheartsuit{}^{\varheartsuit}$$\varheartsuit{}^{\varheartsuit}$epartment of Computer Science, University of Southern California Information Sciences Institute, University of Southern California sramnath@usc.edu
1 Introduction
Figure 1: Cave generates structured free-text explanations that can be automatically verified for consistency.
Authorship Verification (Av) Koppel and Winter (2014) is the NLP task of determining if two input documents were written by the same author. Av is used for tasks such as plagiarism detection, forensic analysis (often in support of law enforcement), analysis of the spread of misinformation. Given the sensitive nature of this task, it is imperative to develop Av methods that are explainable, accessible and secure. Explainability ensures that the system’s decisions can be trusted by user. Accessibility ensures easier downstream usability; for example, a well-structured free-text explanation with clearly defined topic points is more easily understood than unstructured explanations such as long lists of low-level features. Finally, security is necessary to safely use the Av system for proprietary data; a locally-hosted model on the user’s server is more secure than an API call to an online model, which could potentially store the user’s sensitive data. Av is also an extremely difficult task for humans, in contrast to other NLP tasks such as machine translation; this further necessitates the need for automatic Av systems that accessible to the user, accurate and explainable.
Early works performed Av by comparing hand-crafted features such as n-grams van Dam (2013), POS tags Moreau and Vogel (2013), LIWC features Uchendu et al. (2020); while these are explainable, they are hard to scale. Subsequent works employed neural architectures to improve scalability and accuracy, such as the usage of Siamese networks Araujo-Pino et al. (2020); Najafi and Tavan (2022); Boenninghoff et al. (2019) to assign authorship via distance between document embeddings. These methods, however, have little or no post-hoc explainability. This poses a huge problem: one cannot, for instance, make legal assertions about the authorship of a set of questionable documents solely on the basis of “a system said so” - decision makers (judges, university officials, intelligence analysts, etc.) need to know why.
Recent years have seen the surge of large language models and their self-rationalization capabilities. Subsequently, there have been works Hung et al. (2023); Huang et al. (2024) that generate free-text rationales [^2] We use terms explanation and rationale interchangeably. for Av from models such as GPT-4 OpenAI (2024b) in a zero/few-shot manner. These methods however face issues such as unreliability of online models (varying results/details across runs, deprecation of APIs), inconsistency of rationales with predicted label Hung et al. (2023) (refer to Appendix B), and high expenses.
In this work, we propose our model Cave (Controllable Authorship Verification Explanations), an in-house model that generates authorship predictions along with high-quality, free-text rationales that explain the former post-hoc. The rationales are structured as a set of sub-explanations that reason about documents’ shared authorship in terms of Av-relevant linguistic features Hung et al. (2023); Boenninghoff et al. (2019); these sub-explanations have corresponding (intermediate) labels that provide further structure and also serve as a means to verify overall consistency with Cave’s authorship prediction.
Since there is no human-written Av explanation data that can be used to train our in-house model, we adopt the model distillation strategy: West et al. (2022); Li et al. (2023a, 2024) we engineer a prompt to generate rationales in the structure described above, generate silver-standard training data by prompting a large teacher model (GPT-4-Turbo OpenAI (2024a)), and use that to finetune a smaller student model (LLaMa-3-8B Meta-AI (2024)). We also introduce our automatic metric Consistency that measures the consistency between the rationale and the model’s final label; Consistency is used to filter the silver train data before finetuning Ramnath et al. (2023); Li et al. (2023a), and also assess Cave during test time. Our experiments with three difficult Av datasets show that Cave leads to competitive task performances as well as high quality of rationales.
2 Our Method
Cave’s Prompt:
Task: On a scale of 0 to 1, with 0 indicating low confidence and 1 indicating high confidence, please provide a general assessment of the likelihood that Text 1 and Text 2 were written by the same author. Your answer should reflect a moderate level of strictness in scoring. Here are some relevant variables to this problem.
1. punctuation style (e.g. hyphen, brackets, colon, comma, parenthesis, quotation mark)
2. special characters style, capitalization style (e.g. Continuous capitalization, capitalizing certain words)
3. acronyms and abbreviations(e.g. Usage of acronyms such as OMG, Abbreviations without punctuation marks such as Mr Rochester vs. Mr. Rochester,Unusual abbreviations such as def vs. definitely)
4. writing style
5. expressions and Idioms
6. tone and mood
7. sentence structure
8. any other relevant aspect
Provide the answer in a Python JSON format. Use the following keys for your dictionary: ‘punctuation style’, ‘special characters style, capitalization style’, ‘acronyms and abbreviations’, ‘writing style’, ‘expressions and Idioms’, ‘tone and mood’, ‘sentence structure’, ‘any other relevant aspect’, ‘final score’. Apart from the ‘final score’, everything else must have a text value; also, the text should include a concluding YES/NO/MAYBE about whether the two texts are similar or not with respect to the key at hand. Finally, provide an ‘output’ key in your dictionary, which says YES if the two texts are written by the same author, and NO otherwise.
Text1: excellent until a completely unsatisfying ending This movie really got me hooked. The plot about…
Text2: A Nutshell Review: Oliver Twist “Please sir, may I have some more?” Ah, the immortal line from…
Table 1: Cave’s prompt: We extend Hung et al. (2023) to improve controllability in terms of the structure/parsing of the output and automatic verification of the consistency between the rationale and the label. The highlighted part is our extension to PromptAV.
Cave’s Output Format:
punctuation style: Both texts use a variety of punctuation, including commas, periods, and quotation marks, but Text 1 uses more diverse punctuation such as parentheses and hyphens. MAYBE
special characters style, capitalization style: Text 1 uses continuous capitalization for emphasis (e.g., ’WONDERFUL’, ’THRILLED’), which is not observed in Text 2. NO,
acronyms and abbreviations: Neither text makes significant use of acronyms or unusual abbreviations, maintaining a formal tone. Conclusion: YES,
writing style: Text 1 has a more personal, reflective style, sharing personal opinions and feelings about the movie. Text 2 provides a more detached, narrative-style review without personal input. NO,
expressions and idioms: Both texts avoid colloquial expressions and idioms, opting for a more formal review format. YES,
tone and mood: Text 1 has a more varied tone, from enthusiasm to disappointment, while Text 2 maintains a consistent, somewhat formal and analytical tone. NO,
sentence structure: Text 1 features a mix of short and long sentences with more complex structures, while Text 2 tends to use more uniformly structured, intermediate-length sentences. MAYBE,
any other relevant aspect: The approach to movie critique is different; Text 1 is more about the impact on the viewer, while Text 2 focuses on plot summary and cinematic elements. NO,
final score: 0.375,
output: NO
Table 2: Sample output by Cave, with the output in a JSON format. The text (JSON keys) highlighted in blue depict the various linguistic features considered for the analysis, as well as the keys for final confidence score and predicted task output. The text underlined and highlighted in yellow depict the intermediate labels with respect to each linguistic feature, the confidence score value, and the predicted task label. We use the predicted label to measure task Accuracy and all of them to measure explanation-label Consistency.
Our proposed pipeline consists of three parts: (1) generating silver rationales with a desired structure from GPT-4-Turbo via a carefully engineered prompt, ‘Cave’s prompt’, (2) filtering them according to our criteria of output format and Consistency, and (3) distilling these into a small language model to get the final Cave. Figure 2 shows a visual demonstration of the pipeline.
Figure 2: Our pipeline. We sample silver rationales from GPT-4-Turbo using Cave’s prompt (Table 2), filter them according to our desired properties and finally distill them into the much smaller LLaMa-3-8B.
2.1 Cave’s prompt
To design Cave’s prompt, we started off by observing the strengths and weaknesses of prior prompting-based methods for Av explanations, PromptAV Hung et al. (2023), Lip Huang et al. (2024). The main strength of these prompts is their compilation of Av-relevant linguistic required for authorship analysis. The main weakness is the inaccessibility of their generated explanations towards automatic rationale quality evaluation333this is reinforced by these works reporting only task accuracy and no rationale quality metrics. Further, human analyses (Appendix B) of rationales generated by PromptAV and CoT (chain-of-thought Wei et al. (2022), a baseline self-rationalization method also presented in Hung et al. (2023)) show that these rationales can be potentially inconsistent with their final label, rendering them untrustworthy.
We adopted PromptAV as our base prompt, and performed several iterations of prompt engineering on GPT-4-Turbo alongside human analyses of rationale quality and downstream utility. Our final ‘Cave’s prompt’ (Table 2) generates a highly structured output rationale that can be easily decomposed for (1) verifying rationale-label consistency and (2) more accessible layman utility; Cave’s prompt retains all the linguistic features compiled by PromptAV Boenninghoff et al. (2019). As shown in Table 2, Cave’s outputs follow a JSON-type text format where the linguistic features are keys; for each feature, the model has to provide an explanation and an intermediate output (YES/NO/MAYBE) that denotes whether this feature indicates shared authorship. Finally, the model has to provide a score from 00 to $1$ which indicates its confidence that the documents have the same author, and a YES/NO task label output. The output’s JSON format enables us to easily parse the explanation into its different sub-steps in an automatic manner; (re-iterating) this enables the downstream utility and automatic evaluation of the rationale.
2.2 Automatic Evaluation & Filtering
We define the following two criteria to evaluate the structure and consistency of rationales. We use these criteria in the distillation process (Section 2.3) to filter out low-quality silver training data, and at evaluation time.
Structure.
We define the correct and complete output structure to be a JSON format with all the target linguistic features, final score and final label.
Rationale Consistency.
Rationales that are inconsistent with the model’s predicted label are risky and unusable in downstream applications Lyu et al. (2023). Prior works in self-rationalization check for consistency either by human analysis Chen et al. (2023b) or by training models Hase et al. (2020); Chen et al. (2023a); Wiegreffe et al. (2021) that can do the same. In Cave’s output format, the presence of intermediate and final labels allows us to automatically verify rationale-label consistency. We define a Consistency metric using two binary $0/1$ variables Cs-1 and Cs-2:
Cs-1: Is the ‘final score’ (which indicates the model’s confidence in whether the two documents have the same author) consistent with the final label? If the final label is ‘YES’, is the final score $\geq 0.5$? Similarly, if the final label is ‘NO’, is the final score $\leq 0.5$?
Cs-2: Are the intermediate labels faithful to the final label? Since the model output is structured, it is easy to automatically extract the YES/NO/MAYBE labels corresponding to each dimension of analysis and count them. If the final label is ‘YES’ and count(YES) + count (MAYBE) > count(NO), then the labels are consistent. Similarly, if the final label is ‘NO’ and count(NO) + count(MAYBE) > count(YES) then the labels are consistent.
We define $\textsc{Consistency}=\frac{\textsc{Cs-1}+\textsc{Cs-2}}{2}$. That is, a given output is consistent ($\textsc{Consistency}=1$) only if both Cs-1 and Cs-2 are $1$.
2.3 Model Distillation
We use Cave’s prompt ($P$) to generate silver training data from GPT-4-Turbo ($T$); we then filter out the low-quality silver data based on output structure and Consistency. We distill this filtered training data into a smaller language model LLaMa-3-8B ($S$). Formally, given the prompt $P$, teacher model $T$ and student model $S$, we generate silver training data:
$D_{train}=\{(x_{i},y_{i}^{t})\}\sim T(x_{i}|P)$ (1)
where $x_{i}$ represents the input document pairs and $y_{i}^{t}$ represents the explanations created from $T$. We filter $D_{train}$ using the criteria of $\textsc{Consistency}=1$ and conformity to the complete JSON format to get $F_{train}$. We then train $S$ with $F_{train}$ with standard language modeling loss:
$E_{(x,y^{t})\sim F_{train}}S(y^{t}|x)$ (2)
3 Experiments and Results
3.1 Datasets
We perform experiments on three popular Av datasets: IMDb62 Seroussi et al. (2014), Blog-Auth Schler et al. (2006) and FanFiction (via the PAN-2020 Authorship Verification task444pan-2020-task). Our datasets are diverse in terms of their domains (IMDb62- movie reviews, Blog-Auth- website blogs, FanFiction- movie/book-based fanfiction); we carefully select these three datasets in order to test the generalizability of our method.
For IMDb62 we use the test set of 1k text pairs created by Hung et al. (2023). For Blog-Auth, we create our own test set by sampling 1k document pairs from the original blog authorship corpus555blog-authorship-corpus. For FanFiction, we sample 1k document pairs from the PAN-2020 test set (which had 14k pairs).
Creating the training set.
For IMDb62 and Blog-Auth, the original datasets consist of documents and their anonymous author IDs; we sample document pairs with equal amounts of same/different authors to create training datasets for them. For FanFiction, we take the ‘small’ training set provided by PAN-2020, and sample datapoints from the same. We start off with a training set of 1.8k document pairs per dataset. We then combine them with Cave’s prompt and get silver rationale data from GPT-4-Turbo (temperature $0.0$ and $2$ responses per sample); this results in an initial set of 3.6k datapoints. Finally, we filter them according to the criteria in Section 2.2 and use them for training - our final training data sizes are 2.3k, 2.7k and 2.1k for IMDb62, Blog-Auth and FanFiction respectively.
3.2 Training and Evaluation details
We distill the above silver training data into a LLaMa-3-8B, which is orders of magnitude smaller than GPT-4-Turbo. We train it with LoRA Hu et al. (2021), a parameter-efficient training method. We report all hyperparameters in Appendix C. We train models in both single dataset and multi-dataset settings and report metrics Accuracy, Consistency in Table 3.
Model Method Test $\rightarrow$ IMDb62 Blog-Auth FanFiction
Train $\downarrow$ Acc. Cons. Acc. Cons. Acc. Cons.
SVM - Single-dataset 59.7 - 56.2 - 55.7 -
GPT-4-Turbo CoT Zero-shot 73.6 - 62.0 - 54.2 -
PromptAV Zero-shot 73.0 - 62.2 - 57.0 -
Cave’s prompt Zero-shot 67.6 0.86 57.7 0.73 52.2 0.67
Cave’s prompt Few-shot 71.3 0.94 64.0 0.97* 61.7* 0.86
CoT Zero-shot 62.4 - 60.7 - 57.5 -
LLaMa-3-8B PromptAV Zero-shot 62.8 - 57.9 - 53.8 -
-Instruct Cave’s prompt Zero-shot 36.7 0.16 38.8 0.25 45.5 0.17
Cave’s prompt Few-shot 53.8 0.98 62.3 0.99 48.1 0.93
LLaMa-3-8B CoT Single-dataset 62.0 - 64.2* - 61.5 -
Multi-dataset 75.3 - 62.5 - 57.6 -
PromptAV Single-dataset 63.8 - 58.4 - 58.2 -
Multi-dataset 68.6 - 62.7 - 63.3 -
LLaMa-3-8B Cave Single-dataset 75.7* 0.97 60.4 0.94 60.2 0.81
Multi-dataset 74.9 0.99* 57.9 0.96 58.6 0.91*
Table 3: Empirical Results comparing baselines and Cave. We note that both single dataset and multi-dataset trained models beat/are competitive with strong baselines such as GPT-4-Turbo. Note that the single-dataset results each correspond to three different Cave models trained on the respective training sets.
3.3 Baselines
We present and compare with multiple baselines:
Traditional stylometry.
We present a weak baseline Ikae (2021) from traditional stylometry. We represent input document pairs as the concatenation of their TF-IDF666sklearn.feature_extraction.text.TfidfVectorizer or word-count 777sklearn.feature_extraction.text.CountVectorizer features, and train a Support Vector Machine Cortes and Vapnik (1995) to classify their shared authorship status. We report final hyperparameters for the same in Appendix C.
Zero-shot, prior work.
We present three zero-shot baselines with GPT-4-Turbo (temperature $0.0$) and LLaMa-3-8B-Instruct (greedy decoding). The first baseline is chain-of-thought CoT Wei et al. (2022), a standard self-rationalization method that does not include any Av specific instructions. The second baseline is PromptAV Hung et al. (2023). We use the exact prompts for CoT and PromptAV as reported in Hung et al. (2023) (refer Appendix B). Both these baselines report confidence scores (0-1, where higher score means more confidence in shared authorship) in lieu of labels. We extract this confidence score by string matching for decimal numbers in the output; we measure Accuracy by thresholding these scores at $0.5$. Consistency cannot be calculated for these two baselines. The third zero-shot baseline is Cave’s prompt itself.
Finetuned.
Similar to how we distill Cave, we generate silver CoT and PromptAV data from GPT-4-Turbo, filter them according to task correctness and distill them into a LLaMa-3-8B. We report final train set sizes in Appendix C.
Few-shot.
Finally, we present a few-shot baseline with Cave’s prompt on GPT-4-Turbo and LLaMa-3-8B-Instruct. This baseline is strong, but undesirable owing to the high expenses of performing few-shot Av  [^8] Each individual document in our datasets has on-average 1000 tokens; performing even a 4-shot baseline is highly expensive.. We have submitted the few shot prompts in supplementary material.
3.4 Empirical Results
Table 3 shows the empirical comparison between the baselines and Cave. We report Accuracy for all models, and Consistency for Cave.
Single-dataset training.
We first look at the LLaMa-3-8B models trained on individual datasets. For IMDb62, Cave beats all baselines (including zero and few-shot GPT-4-Turbo, which is orders of magnitude larger in parameter size) on both accuracy and consistency; for FanFiction, Cave beats all baselines except for few-shot GPT-4-Turbo. For Blog-Auth, Cave beats significant baselines such as zero-shot GPT-4-Turbo, and is competitive with the remaining baselines in terms of both metrics.
Multi-dataset training.
We also train a multi-dataset Cave on all 3 datasets together to observe any transfer-learning trends (the last row in Table 3). This model gives the best consistency for all datasets. There is minimal loss in accuracy when we move from single dataset to multi-dataset training, which is complemented by increased consistency; for IMDb62, multi-dataset Cave still beats all baselines in terms of both metrics.
Why show two zero/few-shot baselines?
In addition to the strong (online) baseline of GPT-4-Turbo, we also show zero/few-shot results using a less-strong local model, LLaMa-3-8B-Instruct. The significant performance degradation there (for all cases except few-shot Blog-Auth) shows that it is insufficient merely to take existing approaches and run them using a local instruction-tuned model; our proposed distillation strategies are essential to get a high performance with local, offline models that face the disadvantage of having a much lower parameter size.
LLaMa-3-8B for Cave.
In Appendix D, we report results on Cave with other offline student models Jiang et al. (2023); Team (2024); Stability-AI (2024) of varying sizes to inform design choices in downstream applications; we find that LLaMa-3-8B is the best performing Cave, with Mistral-7B Jiang et al. (2023) being the second-best.
3.5 Human Pilot
Prior works in this field largely focus on task accuracy, with much lesser focus on explainability Patel et al. (2023) and by extension, human study of the explanations. In this work, we take the first step towards human analysis of Av explanations by providing a pilot study on Cave’s rationales.
In-house human study: Av + explanations is a highly complex task; annotators require a considerable amount of training and practice to be skillful in their evaluation. Hence, for our evaluation, we opted to conduct a small but high-quality in-house human study, instead of a large-scale study with platforms such as Amazon MTurk.
We pick $20$ document pairs per dataset (total of $60$ pairs) and evaluate the rationales generated by multi-dataset Cave. We analyze the strengths & weaknesses of Cave’s rationales, which help to understand the utility of these rationales to downstream applications Joshi et al. (2023) and to learn about potential areas for improvement van der Lee et al. (2019). We analyze the generated rationales (for each linguistic feature as in Table 2) via three properties as described below; we employ 3 distinct annotators per datapoint. We also require the annotators to comment on why a rationale is unsatisfactory, if they find it to be so with respect to any property; this helps to perform a more detailed analysis at the instance level.
1. Detail-Consistency: Are the details mentioned in the rationale consistent with the input documents, or are they hallucinated? (eg:- if the rationale mentions that both documents use parantheses, but the documents don’t, then the details are hallucinated)
2. Factual-Correctness: Are the rationales factually correct? (eg:- if the rationale says that the text was informal in tone, while it was actually formal, or if it misinterprets an acronym to be author-specific slang, it is incorrect)
3. Label-Consistency Cave’s output format has a distinct rationale and intermediate label (YES/NO/MAYBE) for each linguistic feature. As defined in Section 2.2, we use these intermediate labels to check the overall consistency of the rationale with the final label. In this human evaluation however, we check whether each individual rationale is consistent with its corresponding label.
Linguistic Features P1 P2 P3
special characters & capitalization style 35 37 32
acronyms and abbreviations 47 40 24
writing style 52 54 56
expressions and idioms 51 48 42
tone and mood 55 54 57
sentence structure 49 53 44
any other relevant tabs 53 56 41
aggregate 342 342 296
aggregate as % 81.4% 81.4% 70.5%
Table 4: Human Pilot: This table shows the number of examples out of 60 where all 3 annotators agree that the rationale conforms to the property being analyzed. P1 - ‘Detail-Consistency’, P2 - ‘Factual-Correctness’, P3 - ‘Label-Consistency’. We report these numbers for all features except punctuation (discussed in Section 3.5.
Rationale / Label Annotator’s Comment
punctuation style: Both texts use a variety of punctuation marks such as commas, parentheses, and hyphens consistently. Hallucination of punctuation marks when they don’t exist in the original documents.
any other relevant aspect: The content of both texts revolves around personal experiences, daily life, and interactions with others, suggesting a similarity in thematic focus. YES Similarity/difference in topics does not definitively imply same/different authorship - the label should have been MAYBE for both these cases.
any other relevant aspect: The thematic content of Text1 is significantly different from Text2, with the former focusing on a fictional narrative and the latter on a real-world emotional journey. NO
special characters style, capitalization style: Text1 shows standard capitalization with no special characters. Text2 uses URL links and capitalizes ’RICE’ as an acronym, indicating a different approach to capitalization and special characters. Capitalizing an acronym is not a ‘different approach’ to capitalization.
acronyms and abbreviations: Neither text heavily relies on acronyms or unusual abbreviations, maintaining a formal tone. Lack of acronyms does not indicate a formal tone - this rationale is even more incorrect when the original text is informal in tone.
expressions and idioms: Text1 uses colloquial expressions (’nice???’, ’hehe nice rite’). Text2 uses a known idiom (’your princess is in another castle’) in a humorous context. NO The two expressions/idioms are not conflicting, and hence this isn’t enough a reason to say NO. The label should have been MAYBE.
punctuation style: Both texts use quotation marks inconsistently and have a similar approach to dialogue punctuation, indicating a possible match. Dataset error in FanFiction, where double quotes are used as apostrophes. The dataset error is not an indicator of authorship, but the model doesn’t know that.
Table 5: Sample rationales from various examples, to show observed error types by Cave
In Table 4, we show number of examples (out of 60) for which the three annotators agree that the rationale aligns with the property being analyzed. We also show aggregate statistics for each property across all linguistic features, i.e., $\frac{sum(P1)}{60*7}$ where 7 is the number of linguistic features being analyzed. We report numbers for all linguistic features except for punctuation style, since the FanFiction dataset suffers from a dataset bias error for the same (discussed further below).
Aggregate Analysis. Annotators find that Cave’s rationales align with properties P1 and P2 in 81.4% of the cases, and with property P3 in 70.4% of the cases. That is, Cave on average hallucinates or presents incorrect details in $\sim$19of the cases, and produces inconsistent intermediate labels in $\sim$30of the cases.
Fine-grained analysis. Out of the 21 pairs of linguistic features (7) x properties (3) analyzed, annotators find that $\geq 70\%$ of the datapoints are aligned with the property analyzed for 15 pairs. For the cases where annotators finds the rationales to be dissatisfactory (such as ‘capitalization’, ‘acronym’) we take a deeper look into annotators’ comments.
Errors due to dataset biases. Annotators noticed that documents in the FanFiction dataset had a bias - they all used double instead of single quotes for apostrophes (eg:- it“s instead of it‘s), leading to persistent errors across all examples - Cave concluded that these double quotes were indicative of authorship, and consistently marked that FanFiction document pairs had the same author with respect to punctuation style.
Qualitative examples & Annotator comments. We report commonly observed rationale errors in Table 5. Hallucination of details was a common error, especially for punctuation style and special characters / capitalization; for example, Cave cited the presence of punctuation marks such as ellipses, parantheses, etc. as the similarity/difference between documents, when in reality, those did not exist in the document. One hypothesis is that LLMs are bad at these linguistic features since they deal with individual characters and not broad pieces of text like the other features. Another common error was the Cave’s usage of topic of texts as a reason to indicate or refute shared authorship of two documents; topic cannot be an absolute factor in deciding authorship Wegmann et al. (2022a), and hence such labels must be ‘MAYBE’. There were also cases where Cave used trivial grounds to reason about authorship (for example, ‘documents have the same author because both have standard acronyms and capitalization’, more examples in Table 5); such trivial reasons should give rise to an intermediate label of ‘MAYBE’ and not ‘YES/NO’ - most of the cases where the annotators marked the rationales as unsatisfactory for property P3 were cases where Cave gave a ‘YES/NO’ label and the annotators deemed it should be a ‘MAYBE’.
Initial steps towards handling these issues include reduced reliance on linguistic features where Cave struggles such as punctuation style; this includes (1) making downstream users aware about them so that they can make an informed final decision with Cave’s rationales, (2) modifications to Cave’s prompt so that these features are used to make authorship conclusions only if the model observes unusual behaviour in that regard. As future work, we would like to address these shortcomings with targeted training. We would also like to look at dynamic weighing of linguistic features at inference time, i.e., which features contributes more/less to the authorship analysis of these two documents? Finally, we also discuss random errors owing to computational limitations (such as limited context window in LMs) in Appendix G.
We report these numbers for individual datasets in Table 12 in Appendix F. We also release the complete set of test-set rationales generated by Cave  [^9] Cave’s rationales Sheet.
4 Discussion
4.1 Accuracy vs. Explainability
Early works on self-rationalization such as Kojima et al. (2022); Marasovic et al. (2022) showed that rationale generation improved task accuracies; however, most such works place a higher focus on task performance than on the rationale quality. Recent works Joshi et al. (2023); Ramnath et al. (2023) discuss the need for increased rationale quality in order for them to be useful and trusted in downstream applications. Hence, in our efforts towards better Av, we place a high importance on rationale quality; Cave achieves both increased explainability (via consistency of labels) as well as competitive task accuracies.
We measured consistency. What next?
In addition to consistency, there are several properties necessary for the rationale to be of good quality Joshi et al. (2023); Ramnath et al. (2023). Some of these properties are easily verifiable by humans, even if there exist no prevalent automatic methods to measure them: for eg:- factual correctness, hallucinated details in the rationale, etc. that we measured in our human evaluation (Section 3.5). There are also properties that are hard to verify even by human evaluators: for eg:- completeness of the rationale (i.e., has the rationale missed any significant similarity/dissimilarity between the documents?). In fact, completeness does not have a comprehensive definition even for simpler tasks like multi-choice QA Ramnath et al. (2023). In the future, we would like to work on automatic metrics and model training for these properties in Av.
4.2 Potential Risks
As discussed in Section 1, Av is a high-stakes task. While we strive to enforce explainability in Cave’s generations, it is still a language model which is trained on data generated from a black-box LM; both GPT-4-Turbo’s and LLaMa-3-8B’s biases could have been propagated to Cave. Further, as we see in Section 3.5, Cave’s rationales sometimes have hallucinated details, especially for certain linguistic features. Hence, we caution downstream users that Cave’s outputs should be used only after due understanding and verification. We discuss more details in the Limitations section at the end.
5 Conclusion
Existing methods for authorship verification face a lack of accessible explainability. In this work, we train an in-house language model Cave to generates structured and consistent explanations for Av. We present strong experimental results on three difficult Av datasets; we also present a human pilot study on the quality of generated rationales. Finally, we provide detailed discussions on explainability in Av and explore errors made by Cave due to factors such as hallucination, dataset biases, etc. to guide research in this area. As future work, we would like to explore reward metrics and reward-based learning for Av explanations.
Acknowledgements
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200006. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein. We thank the human annotators for their help with the pilot study. Finally, we thank all our collaborators at USC INK Research Lab and ISI for their constructive feedback on this work.
Limitations
Like any large language model, Cave can unintentionally lead to toxic, incorrect or hallucinated text; it is the responsibility of the user to use the model mindfully. Further, since Cave is designed for the task of Av which is extremely sensitive in nature, it is the responsibility of the user to duly verify the data used to train the model, and also verify the output given by the model before using it. Further, as noted in the main paper, our models have not been explicitly trained for completeness of rationale; that is, apart from wrong information the generated explanation may have, it could also have missed some crucial similarity/dissimilarity between the documents.
Further, the Cave models trained in this paper were trained on data sampled from yet another language model GPT-4-Turbo and not on human-written gold standard data; it is possible that harmful biases present in GPT-4-Turbo (by process of its training and its training data) might have been passed over to the silver training data. It is necessary to keep this in mind while using our released training data.
Lastly we note that our method has practical limitations faced by any NLP task; our models are restricted by the size of the context window allowed by the model (both input and output text) (also keeping in mind restrictions due to computational costs, financial costs and GPU size available). This can lead to data missed by the model in its reasoning.
Data.
All the datasets that we use in our work are released publicly for usage and have been duly attributed to their original authors. Note that while we perform authorship verification, all the datasets we use have been anonymized by their respective creators. We do not have access to or use the names of individual people.
Reproducibility
For all our experiments, we report (1) the complete hyperparameter setting and any bounds explored (Appendix C) as well as the sizes and versions/pretrained-model links of all models used, (2) the time taken per experiment, and infrastructure used, (3) descriptions of datasets, and demonstrations used to sample rationales from GPT-4-Turbo. Further, we have also submitted all codes as supplementary material.
Llama-3 Usage and License.
Note that we were granted access to download and use LLaMa-3-8B from https://huggingface.co/meta-llama/Meta-Llama-3-8B. While we train a local LLaMa-3-8B, we do not release the trained model; we submit only the script and data used to train the model. Also note that we adhere to the intended use case for LLaMa-3-8B, i.e., “Llama 3 is intended for commercial and research use in English. Instruction tuned models are intended for assistant-like chat, whereas pretrained models can be adapted for a variety of natural language generation tasks.”
Appendix A Related Work
A.1 Authorship Analysis
Classical AV systems use inter-textual distance to quantify the likelihood for two documents to share the same author Mosteller and Wallace (1984); Burrows (2002); Grieve (2007). These verification techniques rely on expert-derived hand-crafted features to define the distance metric. Since such stylometric methods are grounded in hand-written rules, they are inherently explainable. However, these methods are known to not scale well, especially in the scenario when for a given query document the correct target author needs to be mined from a large collection of candidate documents [^10] See Table 1 of Rivera-Soto et al. (2021) for a comparison between scores from neural transformer/convolutional models and classical tf-idf technique..
Different from classical stylometric techniques, current AV systems typically employ neural architectures which perform better at scale Rivera-Soto et al. (2021); Wegmann et al. (2022b). Although significant amount of work has been devoted to advancing neural AV architectures, most of the current techniques lack interpretablity. For example, a typical neural AV architecture comprises of a Siamese network Koch (2015), where each document is separately encoded into a vector and the model is trained to ensure that the document vectors are close if the they are written by the same author and far otherwise Rivera-Soto et al. (2021); Wegmann et al. (2022b). In such a network, what each dimension of the document vector means is completely opaque.
To bridge this dichotomy between usability and interpretibilty between classical stylometric techniques and neural AV systems, Boenninghoff et al. (2019) proposed to use attention heat-maps. This technique enabled them to determine which tokens in the input text are attended to (and not attended to) by the neural model when performing AV. Different from attention-heat maps, model-agnostic methods, such as LIME Ribeiro et al. (2016), have also been used to probe feature importance Sari et al. (2018). Such explanations are helpful, however, they are inherently local [^11] For example, heat-maps can only capture features at the token/span level. and thus cannot capture higher level abstractions which can be informative of authorship, for example fluency, tone and mood, etc.
Motivated by the recent advancements made possible using LLMs, Hung et al. (2023) and Huang et al. (2024) investigated whether LLMs can perform the task of authorship analysis. They found that, when prompted correctly, the LLMs can reliably determine whether or not two documents are written by the same author. LLMs have also been used to construct interpretable authorship style vectors Patel et al. (2023) and subsequently used to rewrite text from a source author in the writing style of a target author Patel et al. (2022).
Our work is the first attempt to distill free-text explanations for authorship verification from LLMs like GPT-4-Turbo into smaller LMs; we provide explicit focus to the joint explainability and security of Av systems via our model Cave.
A.2 Self-rationalization
Explainability for neural netwrk based models come in two formats: extractive rationales (includes pre-LLM era) and free-text rationales. Extractive rationales (Li et al., 2016; Sundararajan et al., 2017; Lundberg and Lee, 2017; Jin et al., 2019) involve analyzing the influence of input tokens on the predicted output via various methods such as gradient-based analysis of input tokens  (Sundararajan et al., 2017; Lundberg and Lee, 2017), input perturbation (Poerner et al., 2018; Kádár et al., 2017), attention heatmap analysis  (Pruthi et al., 2020; Stacey et al., 2022; Wiegreffe and Pinter, 2019), and trained models for this purpose (Lei et al., 2016; Chan et al., 2022; Jain et al., 2020; Situ et al., 2021; Liu et al., 2023). However, extractive rationales have limited applicability as discussed previously; hence we focus on free-text rationales.
The initial set of approaches for free-text rationales used gold standard human-written rationales to train rationale generation models (Camburu et al., 2018; Narang et al., 2020; Wiegreffe et al., 2021). Following the advent of emergent self-rationalization in LLMs Wei et al. (2022), the research area moved to prompt large LMs with the help of curated templates with or without rationale demonstrations (Wei et al., 2022; Kojima et al., 2023; Li et al., 2023b; Jung et al., 2022; Lightman et al., 2023). Other approaches include leverage few-shot training approaches with a handful of gold rationales (Marasovic et al., 2022; Chen et al., 2023b), or generating silver standard training data from large LMs to distill small LMs to be better at the task or better rationalizers. (Ramnath et al., 2023; Li et al., 2023a; Chan et al., 2023; Wang et al., 2023; Saha et al., 2023; Hsieh et al., 2023).
Appendix B PromptAV and CoT
CoT:
Task: On a scale of 0 to 1, with 0 indicating low confidence and 1 indicating high confidence, please provide a general assessment of the likelihood that Text 1 and Text 2 were written by the same author. Your answer should reflect a moderate level of strictness in scoring. Let’s think step by step.
Text1: [T1]
Text2: [T2]
Table 6: CoT prompt for the task of Av, taken from Hung et al. (2023)
PromptAV:
Task: On a scale of 0 to 1, with 0 indicating low confidence and 1 indicating high confidence, please provide a general assessment of the likelihood that Text 1 and Text 2 were written by the same author. Your answer should reflect a moderate level of strictness in scoring. Here are some relevant variables to this problem.
Text1: [T1]
Text2: [T2]
Table 7: PromptAV prompt for the task of Av, taken from Hung et al. (2023)
Tables B and B show the prompts used to apply CoT and PromptAV respectively. The prompts are taken from Hung et al. (2023); further note that the prompt is uniform across datasets.
Sample CoT explanation that is inconsistent with label:
Based on the provided texts, I assess the likelihood that Text 1 and Text 2 were written by the same author as 0.7 (moderate confidence).
Here’s my reasoning:
1. Both texts exhibit a personal and introspective tone, which suggests that the author is comfortable sharing their thoughts and feelings.
2. The language and sentence structure in both texts are similar, with a mix of short and long sentences, and a conversational tone.
3. Both texts touch on themes of self-reflection, personal growth, and the struggle to overcome negative thoughts and emotions.
4. The author’s writing style is informal and conversational, with a focus on expressing their thoughts and feelings rather than presenting a polished or formal argument.
However, there are some differences that suggest the texts may not have been written by the same author:
1. The tone of Text 1 is more lighthearted and humorous, while Text 2 is more serious and introspective.
2. The language and vocabulary used in Text 2 are more formal and philosophical, with words like ‘suffering,’ ‘unrealistic standards,’ and ‘setbacks.’
3. The structure and organization of the texts are different, with Text 1 consisting of short, disconnected paragraphs, and Text 2 featuring longer, more cohesive paragraphs.
Overall, while there are some similarities between the two texts, the differences in tone, language, and structure suggest that they may have been written by different authors.
Table 8: Sample output by CoT where the predicted label and the predicted confidence score are inconsistent with each other. The model predicts a high confidence score of 0.7 for same authorship (by usual standards of 0.5 as the threshold), but it predicts that the final answer is different authorship.
Sample PromptAV explanation that is inconsistent with label:
To assess the likelihood that Text 1 and Text 2 were written by the same author, we will analyze the texts based on the provided variables:
1. Punctuation Style:
- Text 1 uses a variety of punctuation, including colons, commas, and quotation marks, with a distinctive style of spacing before and after punctuation marks (e.g., ‘this : THIS’). It also uses apostrophes in plural forms, which is unconventional (e.g., ‘one’s’).
- Text 2 uses standard punctuation without the unconventional spacing or misuse of apostrophes seen in Text 1.
2. Special Characters and Capitalization Style: - Text 1 features continuous capitalization for emphasis (e.g., ‘THIS IS BETTER THAN THE ORIGINAL’) and capitalizes certain words for emphasis.
- Text 2 does not exhibit continuous capitalization or the same pattern of capitalizing certain words for emphasis.
3. Acronyms and Abbreviations:
- Text 1 does not prominently feature acronyms or abbreviations.
- Text 2 also does not prominently feature acronyms or abbreviations, making this variable less useful for differentiation.
4. Writing Style:
- Text 1 has a detailed, narrative style, sharing personal experiences and opinions in depth.
- Text 2 has a more straightforward, opinion-based style without the narrative depth or personal anecdotes found in Text 1.
5. Expressions and Idioms:
- Text 1 uses expressions and a more vivid language (e.g., ‘take 15 showers afterwards’).
- Text 2 uses simpler language and fewer idiomatic expressions.
6. Tone and Mood:
- Text 1 has a passionate, intense tone, especially when discussing the movie’s impact and the author’s feelings.
- Text 2 has a lighter, more casual tone, even when expressing disappointment.
7. Sentence Structure:
- Text 1 features complex sentences with multiple clauses and detailed descriptions.
- Text 2 uses simpler, more straightforward sentence structures.
8. Other Aspects:
- The subject matter and interest areas differ significantly between the two texts, with Text 1 focusing on horror films and Text 2 on a Disney film. This could indicate different interests or purposes for writing, though it’s not definitive evidence of different authorship.
Given these observations, the differences in punctuation style, capitalization, writing style, tone, and sentence structure suggest that Text 1 and Text 2 were likely written by different authors. However, without more direct evidence (e.g., consistent unique spelling errors or very distinctive stylistic quirks that appear in both texts), absolute certainty is not possible.
Confidence Score: 0.8
This score reflects a high level of confidence that the texts were written by different authors, based on the analysis of the provided variables, but it leaves room for the possibility of the same author adapting their style to different subjects or formats."
Table 9: Sample output by PromptAV where the predicted label in the text and the predicted confidence score are inconsistent with each other. The model predicts a high confidence score of 0.8 (by usual standards of 0.5 as the threshold), but it predicts that the final answer is different authorship. Unless a human (or a trained model) can manually analyze the text, it is impossible to determine that the predicted label via the confidence score and the actual predicted label are different.
PromptAV (and by extension, CoT) which serve as the inspiration for Cave produce outputs which cannot be automatically checked for consistency; these prompts produces a free-text explanation followed by a final score which can be thresholded to produce a YES/NO Av label. As observed in the experiments of Hung et al. (2023), models tend to produce low confidence scores with PromptAV, which means that for better accuracy, the optimal threshold in their experiments with GPT-3.5-Turbo was found to be 0.2-0.3, which will most probably not be consistent with the free text rationale (that is, if the rationale says that it thinks it is not confident that the authors are the same, and the confidence is 0.25, the automatic label prediction will still assume that the predicted label is YES). Further, with just the score as the label, it is still possible that we misconstrue the label because the confidence score and the model’s predicted label in the text are different, as shown in the examples in Tables 8 and 9.
Appendix C Hyperparameters
We download the pre-trained LLaMa-3-8B from HuggingFace (https://huggingface.co/meta-llama/Meta-Llama-3-8B) and train it for all our experiments. We train it with Low-Rank Adaption (LoRA; Hu et al., 2021), a parameter-efficient training method which offers lightweight training no additional inference latency, unlike other existing parameter-efficient methods Houlsby et al. (2019). For single-dataset Cave, we train LLaMa-3-8B for $10$ epochs; for multi-dataset Cave, we train for $5$ epochs. We use no validation set, and instead use the last checkpoint obtained from training; since we have no gold-standard explanations, we opt against using the GPT-4-Turbo generated data as the validation set. We experiment with different values of ‘r’ and ‘alpha’ for LoRA (refer to https://huggingface.co/docs/peft/main/en/conceptual_guides/lora for their definitions). For IMDb62 and FanFiction, we use r=128 and alpha=256; for Blog-Auth and for the multi-dataset case, we use r=64 and alpha=128. Note that all our training experiments for LLaMa-3-8B take $\sim$12 hours, and all our test set inferences take $\sim$7-8 hours on a single NVIDIA Quadro RTX 8000 GPU. We report single-run results for all our experiments.
For our baseline SVM, we use TF-IDF feature vectors for Blog-Auth (words) and FanFiction (characters), and word-count features for IMDb62 (characters). All 3 datasets’ models use the RBF kernel. IMDb62 and FanFiction use n-grams of minimum size 1 and maximum size 2, whereas Blog-Auth uses only n-grams os size 2. IMDb62 uses a C-value of 5.0. We used sklearn version 1.2.2 for this baseline.
For our baseline of LLaMa-3-8B finetuned with CoT and PromptAV, our final training dataset sizes are: for IMDb62, Blog-Auth and FanFiction, the CoT sizes - 2.7k, 2.9k, 2.4k and the PromptAV sizes - 2.6k, 3k, 2.6k.
Human Annotation Instructions:
Given two documents, the task is to say whether they are written by the same author (label: YES) or not (label: NO). The model produces a set of explanations (corresponding to different aspects like punctuation, writing style, etc.) for the same each with an intermediate YES/NO/MAYBE label, and finally a YES/NO output label for the task.
The human evaluation looks at the three properties described below, and gives a score of 1, 0.5, 0 or -1 for each of the individual explanations. The "Sample human eval" section gives an example evaluation. If you give a 0.5 or a 0, please write a short explanation what was wrong with the rationale!
Criteria Definition: Options
consistency with details Is the detail consistent with the documents or hallucinated? (for example, explanation mentions parantheses when it doesn’t exist in input document) 1 - all details are consistent, 0.5 - some details are consistent, some are hallucinated, 0 - all details are hallucinated, -1 - I don’t know
factual correctness Are the details factually correct? (for example, it mentions serious writing style when writing style is actually humorous) 1 - all details are factual, 0.5 - some details are factual, 0 - no details are factual, -1 - I don’t know
consistency with predicted label Is the statement faithful with the YES/NO/MAYBE at the end? 1 - yes it is faithful, 0.5 - some details are faithful, 0 - no it is not faithful, -1 - I don’t know
Table 10: Instructions given to annotators for the human pilot study. Sample human evaluations are in Table 11.
Sample evaluation:
Text1: One last update. I’ve decided to wear my khaki pants all day. I get off work at 4:30. I see no need to rush to the 5:05 train. So I’ll take my time and go on the 5:35. I just blowdried my hair and I think it’s going to be okay. I need to eat something. I’ve eaten my two favorite foods, of late, for breakfast: pop tarts and pears. I think it’s going to be rice and peas for lunch.
Text2: Hey guys! I’d like to tell you all that RICE is organising an event, ’Where is the love’. Go to urlLink my site for more info, or go to the urlLink RICE site .It’s going to be great! there will be around 800 people there, and this event will be about how the world’s gone wrong with terrorism, war… etc. Date: 10 July 2004 Place: Willoughby Civic Centre Time: 7pm Ask Ben, Adro or me for more info. We’d like All of you to come! urlLink
GOLD LABEL: NO
Rationale P1 P2 P3
punctuation style: Both texts use a variety of punctuation marks such as periods, commas, and ellipses. However, Text 1 uses more complex structures like colons and parentheses, which are less prevalent in Text 2. MAYBE 0.5 (ellipses / parantheses hallucinated) 0.5 (colons in both) 1
special characters style, capitalization style:Text 1 shows a more natural capitalization style without unusual emphasis on continuous capitalization or certain words. Text 2 capitalizes ’RICE’ and uses all caps for ’URL’, indicating a different approach. NO 0.5 (url not capitalized in Text2) 0.5 (RICE is implied as an organization, it it not a reflection of the capitalization approach) 1
acronyms and abbreviations: Text 1 does not prominently feature acronyms or abbreviations, while Text 2 includes ’URL’ as an abbreviation. This suggests a difference in the use of abbreviations. Conclusion: NO 1 1 1
writing style: Text 1 has a more personal, reflective style, discussing personal routines and thoughts. Text 2 adopts a more informative and promotional tone, inviting people to an event. The styles are quite distinct. NO 1 1 1
expressions and idioms: Neither text heavily relies on idiomatic expressions, but the context in which they write differs significantly, with Text 1 focusing on personal narrative and Text 2 on event promotion. NO, 1 1 1
tone and mood: The tone of Text 1 is casual and somewhat contemplative, while Text 2 is more enthusiastic and outward-facing, aiming to promote an event. These differences suggest different authors. NO 1 1 1
sentence structure: Text 1 uses a mix of short and long sentences, with a more conversational flow. Text 2 also uses a mix but is more structured, likely due to its promotional nature. The sentence structures show some similarities but also differences in context and purpose. MAYBE 1 1 1
any other relevant aspect: TThe subject matter of the two texts is vastly different, with Text 1 focusing on personal habits and reflections, and Text 2 on organizing an event and expressing concerns about global issues. This difference in content and focus further supports the likelihood of different authors. NO 1 0.5 (difference in content doesn’t mean different authors) 0 (should be MAYBE. Difference in content is not difference in authorship)
Table 11: Sample human evaluation shown to human annotators
Linguistic Features IMDb62 Blog-Auth FanFiction
P1 P2 P3 P1 P2 P3 P1 P2 P3
punctuation style 0 1 15 9 15 19 - - -
special characters/ 9 11 6 14 13 17 12 13 9
capitalization style
acronyms/abbr. 14 11 4 15 14 17 18 15 3
writing style 18 18 19 20 19 20 14 17 17
expressions/idioms 16 15 13 18 16 18 17 17 11
tone and mood 19 17 19 19 20 20 17 17 18
sentence structure 18 18 14 15 19 20 16 16 10
any other 18 17 16 18 20 13 17 19 12
Table 12: Human Pilot for individual datasets: This table shows the number of examples (out of 20) where all 3 annotators agree that the rationale (corresponding to the linguistic feature) conforms to the property being analyzed. P1 is ‘Detail-Consistency’, P2 is ‘Factual-Correctness’ and P3 is ‘Label-Consistency’.
Appendix D Cave with different offline models
Model Test $\rightarrow$ IMDb62 Blog-Auth FanFiction
Train $\downarrow$ Acc. Cons. Acc. Cons. Acc. Cons.
LLaMa-3-8B Single-dataset 75.7* 0.97 60.4 0.94 60.2 0.81
Multi-dataset 74.9 0.99* 57.9 0.96 58.6 0.91*
Mistral-7B Single-dataset 66.0 0.91 58.0 0.87 53.6 0.96
Multi-dataset 74.3 0.96 59.6 0.92 57.4 0.83
Stable-LM-Zephyr-3B Single-dataset 42.5 0.67 47.7 0.81 48.3 0.66
Multi-dataset 34.9 0.47 32.7 0.51 17.5 0.22
Gemma-2-2B Single-dataset 24.5 0.36 32.3 0.55 27.9 0.42
Multi-dataset 53.9 0.69 36.9 0.58 52.2 0.77
Table 13: Empirical Results comparing Cave trained with local models of different sizes. Mistral-7B is the next best performing model after LLaMa-3-8B. Gemma-2-2B and Stable-LM-Zephyr-3B however have lower performances; further, the former’s performance increases with multi-dataset training while the latter’s decreases.
We report experiments with local models of varying sizes in Table 13. After LLaMa-3-8B, Mistral-7B is the next best performing model. The smaller models Gemma-2-2B and Stable-LM-Zephyr-3B however have lower performances; further, the former’s performance increases with multi-dataset training while the latter’s decreases.
Appendix E Uncertainty of Cave
Our metric Consistency serves the dual purpose of measuring consistency as well as uncertainty of the model in its prediction. We mark that a rationale’s $\textsc{Consistency}=1$ if and only if the aggregated intermediate labels, final score and final label all match. If they don’t all match, we deem the output as inconsistent and hence uncertain. Hence, the consistency numbers in Table 3 reflect the uncertainty of the model in its output. Specifically, we note that Cave’s outputs show high consistencies (the highlighted/last two rows in Table 3) indicating that Cave is highly certain of its outputs.
Appendix F Human Evaluation
Table B shows the instructions given to the annotators regarding the properties to be analyzed and Table 11 shows a sample human evaluation. We perform an in-house human evaluation, and annotators used were proficient in English and well aware of the high complexity of the task.
Table 12 shows the human pilot study numbers for individual datasets. The aggregate of these numbers were used for the overall analysis in Section 3.5.
Appendix G Random Errors due to Computational Limitations
We acknowledge random errors (i.e., not systematic) that could have occurred in our results owing to limitations of language models and computational resources. Our method uses a language model to read two input documents and provide an output; however, in practice, language models have a fixed maximum length of input due to model limitations as well as GPU memory limitations. This means that if the two input documents together (along with the instructions) do not fit into the input, it becomes necessary to truncate them, which leads to loss of information of their latter parts. This issue is exacerbated when we move towards using in-house models on user servers (which typically have a shorter context window than API-based models like GPT-4-Turbo). In our experiments with LLaMa-3-8B, we used just 1 GPU (to have a realistic user environment), which means that while training we had to restrict the input to be of maximum length 768768768768121212note: during inference, we did not limit the size of the input, but this could pose a problem if the documents became much longer, which could have affected the training for the longer documents. Further, even for generating the training data from GPT-4-Turbo for FanFiction, the excessive size of the documents made the GPT-4-Turbo generations economically infeasible; we cut our the input to to a maximum of 300 words per document, which could have created errors/incomplete rationales that eventually propagated to our downstream training of LLaMa-3-8B.
Appendix H Out-of-domain inference
Test $\rightarrow$ IMDb62 Blog-Auth FanFiction
Train $\downarrow$ Acc. Cons. Acc. Cons. Acc. Cons.
IMDb62 57.7 0.95 58.7 0.92
Blog-Auth 73.8 0.93 50.6 0.74
FanFiction 68.6 0.70 57.4 0.83
Table 14: Out of domain performance by single-dataset Cave models
Finally, we evaluate the single dataset Cave models on the other datasets’ test sets (refer Table 14); we notice robust out-of-domain performance. Notably, the IMDb62 Cave model performs extremely well across the board, with competitive accuracies and consistencies for all datasets. This is a significant result, as it shows the generalizability of this approach to unseen Av datasets: although we use fine-tuning to improve performance over a zero-shot approach, it is not fine-tuning specifically to the domain that is most critical but rather fine-tuning to the structured format of Cave’s output.